Circular No. 5796

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 FW No observations seem to have been made of this interesting new slow-

moving object (cf. IAUC 5730) during the April dark run. The object was detected in three CCD images obtained on May 17 by O. Hainaut with the 1.54-m Danish reflector at the European Southern Observatory, and the following measurements are by R. M. West:

1993 UT  α_{2000}  δ_{2000}  V
May 17.03238  12^{h}23^{m}59.^{s}43  −2^{o}38′27′′.7  23
17.05998  12 23 59.36  −2 38 27.3
17.09281  12 23 59.27  −2 38 26.7

B. G. Marsden, Center for Astrophysics, remarks that these observations constrain the geocentric distance to the range 39-48 AU, and that observations at both the June and July dark runs will be necessary for establishing the orbit with any reliability. He provides the following improved circular elements:

Epoch = 1993 May 13.0 TT  u = 359.^{o}676
Ω = 187.902 }2000.0
a  = 42.340 AU  i = 7.925
1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
May 3 12^{h}24.^{m}69 − 2^{o}43.^{′}5 41.519 42.340 144.^{o}1 0.^{o}8 23.4
13 12 24.17 − 2 39.8 41.628 42.340 134.3 1.0 23.4
23 12 23.76 − 2 36.8 41.758 42.340 124.5 1.1 23.4
June 2 12 23.47 − 2 34.6 41.904 42.340 114.8 1.2 23.4
12 12 23.31 − 2 33.2 42.062 42.340 105.2 1.3 23.4
22 12 23.29 − 2 32.8 42.227 42.340  95.7 1.4 23.4
July 2 12 23.42 − 2 33.2 42.396 42.340  86.2 1.4 23.5
12 12 23.68 − 2 34.6 42.562 42.340  76.7 1.3 23.5
22 12 24.08 − 2 36.8 42.722 42.340  67.3 1.3 23.5
Aug. 1 12 24.60 − 2 39.9 42.871 42.340  57.8 1.2 23.5

SUPERNOVA 1993J IN NGC 3031 Photometry by H. Mikuž, B. Dintinjana, and T. Zwitter, Ljubljana,

Slovenia (comparison stars from Corwin, IAUC 5742): May 16.87 UT, V = 12.39 ± 0.02, B = 13.47 ± 0.03, R = 11.65 ± 0.02, I = 11.21 ± 0.03.

1993 May 18 Daniel W. E. Green